# Combining straight-line and map-based distances to investigate the connection between proximity to healthy foods and disease

### v1

@@ 4.3 Health in the Piedmont Triad
There is between-neighborhood variability in the prevalences of diagnosed diabetes and obesity outcomes seen across the Piedmont Triad (Figure 2). Contrasts are seen between (i) rural and urban neighborhoods but also (ii) neighborhoods surrounding the same major city. Overall, the regional average prevalences for obesity (median $=0.34$, IQR $=[0.31,0.38]$) and diabetes (median $=0.11$, IQR $=[0.09,0.13]$) were comparable to the state averages ($0.1$ and $0.34$, respectively). Most larger, rural neighborhoods had prevalences of diabetes and obesity that were above the state averages. In contrast, many smaller urban neighborhoods saw lower burdens of these diseases. When zooming in on Forsyth or Guilford County, there seem to exist divisions between neighborhoods with low and high prevalences on opposite sides of their major cities’ downtown areas (Supplemental Figure S7). Socioeconomic differences could potentially drive these divisions, further highlighting the complexity in accounting for the complete picture when analyzing social determinants of health.
Figure 2: Choropleth maps of the crude prevalence of diagnosed diabetes and obesity for census tracts in the Piedmont Triad, North Carolina. The gradient for each map is centered at the state median. Data were taken from the 2022 PLACES dataset (Centers for Disease Control and Prevention, 2022). One census tract was excluded because it had a population of zero.
@@ 4.4 Proximity to Healthy Foods in the Piedmont Triad
From the fully queried data ($N=387$ neighborhoods), the following could be learned about straight-line and map-based food access, as well as their relationship. Across all neighborhoods in the Piedmont Triad, proximity to healthy foods based on map-based distances was farther, on average, than the same measure based on straight-line distances, as expected (median proximity $=1.01$ versus $1.51$ miles). The variability of map-based proximity was larger than that of straight-line proximity, also (IQR $=[0.89,2.99]$ versus $[0.57,2.12]$). Still, the two proximity measures were very highly correlated ($R=0.97$), and the relationship between them was linear (Figure 3).
Figure 3: Scatter plot of straight-line versus map-based proximity to healthy food store for neighborhoods (census tracts) in the Piedmont Triad, North Carolina using the fully queried data ($N=387$). The solid line follows the fitted least-squares linear regression fit between $X$ and $X^{*}$, while the dashed line denotes the hypothetical $X=X^{*}$ if there had been no errors in $X^{*}$.
When comparing food access between neighborhoods, the straight-line and map-based measures depicted similar trends across the Piedmont Triad, with better proximity (shorter distances) around urban areas and worse proximity (larger distances) in the rural surrounds (Figure 1). Some rural tracts with poor straight-line proximity had even worse access to healthy foods according to the map-based measure. For example, the nearest healthy food retailer to the population center of Census Tract 9301 within Caswell County was $10.8$ miles away according to the straight-line distances (the worst in the Piedmont Triad). Using map-based distances instead, proximity to healthy foods in this tract was $13.1$ miles (the third-worst instead). There were also instances where straight-line proximity was pretty poor, but map-based proximity uncovered noticeably worse access to healthy foods. For example, Census Tract 9603 within Montgomery County had straight-line proximity of $7.3$ miles (among the 15 worst in the Piedmont Triad) but map-based proximity of $15.8$ miles (the worst in the area).
@@
Next, a multiplicative error mechanism was considered. On average, straight-line proximity underestimated map-based proximity by $30\%$ (i.e., $\hat{\mu}_{W}=0.70$). The distribution of the multiplicative errors was much more symmetric (Supplemental Figure S8). There were still some neighborhoods with much more extreme errors than others; interestingly, they were different from the neighborhoods with the worst additive ones. The most severe multiplicative error ($W=0.063$) was seen for Census Tract 315.04 within Randolph County, among the neighborhoods with additive errors under one mile in magnitude ($U=-0.92$). The standard deviation of the multiplicative errors $\hat{\sigma}_{W}=0.15$ was chosen as the moderate setting for simulations (Section 3.3).
@@ 4.5 Imputation Models for Straight-Line Proximity
For each health outcome, the imputation of map-based proximity to healthy foods $X$ began by fitting the imputation model described in Section 2.3. For diagnosed diabetes, $\hat{\mu}=1.227+1.369X^{*}-0.196\log(Y)\text{ and }\hat{\sigma}=1.057$. For obesity, $\hat{\mu}=3.330+1.372X^{*}-0.318\log(Y)\text{ and }\hat{\sigma}=1.053$. For simplicity and given the high $R^{2}\approx 0.91$, additional covariates ${\mathbf{Z}}$ (e.g., a rural/urban indicator) were not included in the imputation models, but doing so would be straightforward.
Using these estimates, each census tract missing $X$ in the partially queried dataset was multiply imputed $B=20$ times for the imputation analyses. Imputing the missing $X$ values for the unqueried neighborhoods allowed them to contribute to the disease models in Section 4.6. The imputed $X$ values can also be used to supplement the subset of queried data and create a more accurate and complete description of the landscape of food access in the Piedmont Triad than the naive and complete case analyses, respectively. In Figure 1, the map of proximity to healthy foods based on a single imputation (not including $\log(Y)$) is (i) far more comprehensive than the one based on the complete case (ignoring data for $339$ of $387$ census tracts) and (ii) more similar to the gold standard than the one based on the naive.
@@ 4.6 Modeling the Connection Between Proximity and Disease
Finally, the connections between proximity to healthy foods and the prevalence of each health outcome across neighborhoods (census tracts) in the Piedmont Triad were modeled using Poisson regression. In contrast to the simulations, neighborhoods closer together in space could be more similar than to neighborhoods farther away. To evaluate this potential spatial correlation, Moran’s I was calculated for both diseases of interest using the spdep package in R (Pebesma and Bivand, 2023). Specifically, Moran’s I was calculated for the residuals $e_{i}^{*}=Y_{i}-\widehat{Y}_{i}^{*}$ from the naive analysis, where $Y_{i}$ and $\widehat{Y}_{i}^{*}=\exp(\hat{\beta}_{0}^{*}+\hat{\beta}_{1}^{*}X_{i})Pop_{i}$ denote the observed and predicted number of cases in neighborhood $i$, respectively. The parameters $(\hat{\beta}_{0}^{*},\hat{\beta}_{1}^{*})$ were taken from the naive models, which would have fully available data in practice and should be highly correlated with the gold standard ones. Both diagnosed diabetes ($I=0.057$, $p=0.023$) and obesity ($I=0.059$, $p=0.018$) had spatial autocorrelation that would be considered statistically significant at the $\alpha=0.05$ significance level.
@@
A conditional autoregressive (CAR) model was used to capture the connections between proximity to healthy foods and health outcomes while allowing for spatial autocorrelation between neighborhoods that share a border. This model can be defined as a mixed-effects Poisson regression with a random intercept for each neighborhood based on its bordering neighborhoods. Thus, instead of Equation (1),
$\log\{\textrm{E}_{{\boldsymbol{\beta}}}(Y|X)\}=\beta_{0}+\beta_{1}X+r+\log(Pop),$ (2)
@@
was of interest in the Piedmont Triad data, where $r$ denotes a neighborhood-specific random intercept. The vector of all neighborhoods’ random intercepts $\boldsymbol{r}=(r_{1},\dots,r_{N})^{\top}$ is assumed to follow a multivariate normal distribution with mean vector $\boldsymbol{0}$ and covariance matrix $\boldsymbol{\Sigma}$ defined to allow bordering neighborhoods to be correlated. See Waller and Gotway (2004) for more technical details on the CAR model. While the random intercepts $\boldsymbol{r}$ may be correlated with each other, they are independent of the access variables $\boldsymbol{X}=(X_{1},\dots,X_{N})^{\top}$. Using the fitme function from the spaMM package in R (Rousset and Ferdy, 2014), Equation (2) was estimated following the gold standard, naive, complete case, and imputation analysis approaches. For reference, the non-spatial models, estimating Equation (1), were also fit. Additional covariates ${\mathbf{Z}}$ were not included since many of the individual- and area-level factors that might be controlled for here could have already been used in the small area estimation of $Y$ (Centers for Disease Control and Prevention, 2023).
@@ 4.7 Findings from the Piedmont Triad
The prevalence ratio between two census tracts whose proximities to healthy foods differed by $1$ mile ($PR=\exp(\beta_{1})$) was of primary interest. Forest plots of the estimated prevalence ratios ($\widehat{PR}=\exp(\hat{\beta}_{1})$) from each analysis method, along with their 95% confidence intervals (95% CI), are included in Figure 4. Based on the gold standard analysis, worse access to healthy foods (i.e., larger distances $X$ to the nearest healthy foods retailer) was associated with slightly lower prevalence of obesity ($\widehat{PR}=0.994$, 95% CI: $0.989,1.000$). That is, for every $1$ mile farther a census tract’s nearest healthy foods retailer is from its population center, its prevalence of obesity is expected to decrease by $0.6\%$. This result aligns with the previous findings of Kanchi et al. (2021). However, the confidence interval did touch the null ($PR=1$). For diagnosed diabetes, the estimated prevalence ratio was close to the null ($\widehat{PR}=0.997$), and the confidence interval contained one (95% CI: $0.987,1.008$).
@@
Figure 4: Estimated prevalence ratios (with 95% confidence intervals) for proximity to healthy foods and health outcomes in the Piedmont Triad, North Carolina using different analysis methods. Within each health outcome and method, estimates on the right came from the mixed effects model allowing for spatial autocorrelation between neighboring census tracts; estimates on the left came from the model assuming independence between tracts.
For diagnosed diabetes, the naive estimate was biased toward the null, while for obesity, the naive estimate closely resembled the gold standard. The confidence intervals were slightly wider, leading the 95% CI for obesity to cross the null ($\widehat{PR}=0.994$, 95% CI: $0.987,1.001$). Still, the naive analysis led to the same conclusion about the absence of an association between proximity and diagnosed diabetes ($\widehat{PR}=0.999$, 95% CI: $0.985,1.014$).
@@
The estimated prevalence ratio for diagnosed diabetes ($\widehat{PR}=0.986$) from the complete case analysis was farther from the null than that for the gold standard analysis. Still, its wider confidence interval (95% CI: $0.9721,1.002$) captured the null. Notably, the complete case analysis should be asymptotically consistent since $X$ is missing at random (MAR) (Little, 1992) for the unqueried neighborhoods, but there are no such guarantees in a small sample like this one ($n=48$). These differences could also be explained by the higher estimated baseline prevalence, i.e., the predicted disease prevalence in a census tract where proximity to healthy foods was 00 miles (Supplemental Figure S9). The complete-case estimate for obesity ($\widehat{PR}=0.993$, 95% CI: $0.984,1.002$) was in line with the gold standard and naive. Thus, based on the complete case analysis, the conclusion would be that neither health outcome was significantly associated with proximity to healthy foods.

### v2

@@ 4.3 Health in the Piedmont Triad
There was between-neighborhood variability in the prevalences of diagnosed diabetes and obesity outcomes across the Piedmont Triad (Figure 2). Contrasts were seen between (i) rural and urban neighborhoods but also (ii) neighborhoods surrounding the same major city. Overall, the regional average prevalences for diagnosed diabetes (median $=0.11$, IQR $=[0.09,0.13]$) and obesity (median $=0.34$, IQR $=[0.31,0.38]$) were comparable to the state average prevalences ($0.1$ and $0.34$, respectively). Most larger, rural neighborhoods had prevalences of diabetes and obesity above the state averages. In contrast, many smaller urban neighborhoods saw lower burdens of these diseases.
Figure 2: Choropleth maps of the crude prevalence of diagnosed diabetes and obesity for census tracts in the Piedmont Triad, North Carolina. The gradient for each map is centered at the state average (median). Data were taken from the 2022 PLACES dataset. [57] One census tract was excluded because it had a population of zero.
When zooming in on Forsyth or Guilford County, there seem to exist divisions between neighborhoods with low and high prevalences on opposite sides of their major cities’ downtown areas (Supplemental Figure S9). Socioeconomic differences could potentially drive these divisions, further highlighting the complexity of accounting for the complete picture when analyzing social determinants of health. Toward capturing some of these differences, we incorporated an indicator (denoted by $Z$) of neighborhoods being “metropolitan” areas (i.e., having RUCA codes $1$, $2$, or $3$). By this definition, there were $326$ ($84\%$) metropolitan neighborhoods in the Piedmont Triad. The analyses that follow (i) incorporated metropolitan status into imputing proximity to healthy foods (Section 4.5) and (ii) adjusted for metropolitan status directly and allowed for an interaction with proximity to healthy foods in the disease prevalence models (Section 4.6).
@@ 4.4 Proximity to healthy foods in the Piedmont Triad
From the fully queried data ($N=387$ neighborhoods), the following could be learned about straight-line and map-based food access, as well as the relationship between them. Across all neighborhoods in the Piedmont Triad, proximity to healthy foods based on map-based distances was farther, on average, than the same measure based on straight-line distances, as expected (median proximity $=1.51$ versus $1.01$ miles). The variability of map-based proximity was larger than that of straight-line proximity, also (IQR $=[0.89,2.99]$ versus $[0.57,2.12]$). Still, the two proximity measures were very highly correlated ($R=0.97$), and the relationship between them was linear (Figure 3). Importantly, the relationship between $X$ and $X^{*}$ in the fully queried and partially queried datasets was the same, with both datasets agreeing that map-based proximity $X$ was worse (i.e., larger) than straight-line proximity $X^{*}$, particularly at higher values of $X^{*}$. Thus, the partially queried data used to build the imputation model represented the fully queried (gold standard) data well. The same can be said for this relationship after further breaking down by metropolitan versus non-metropolitan neighborhoods (Supplemental Figure S10), except that among the non-metropolitan neighborhoods, the partially queried relationship was slightly steeper (indicating more severe errors expected).
Figure 3: Scatter plot of straight-line versus map-based proximity to healthy food store for neighborhoods (census tracts) in the Piedmont Triad, North Carolina using the fully queried data ($N=387$) or the partially queried data ($n=48$). The solid line follows the fitted least-squares linear regression between $X$ and $X^{*}$, while the dashed line denotes the hypothetical $X=X^{*}$ if there had been no errors in $X^{*}$ (i.e., the line of equality).
When comparing food access between neighborhoods, the straight-line and map-based measures depicted similar trends across the Piedmont Triad, with better proximity (shorter distances) around urban areas and worse proximity (larger distances) in the rural surrounds (Figure 1). Some rural tracts with poor straight-line proximity had even worse access to healthy foods, according to the map-based measure. For example, the nearest healthy food store to the population center of Census Tract 9301 within Caswell County was $10.8$ miles away according to the straight-line distances (the worst in the Piedmont Triad). Using map-based distances, proximity to healthy foods in this neighborhood was $13.1$ miles (the third-worst instead). There were also instances where straight-line proximity was pretty poor, but map-based proximity uncovered noticeably worse access to healthy foods. For example, Census Tract 9603 within Montgomery County had straight-line proximity of $7.3$ miles (among the 15 worst in the Piedmont Triad) but map-based proximity of $15.8$ miles (the worst in the region).
@@
Next, a multiplicative error mechanism was considered. On average, straight-line proximity underestimated map-based proximity by $30\%$ (i.e., $\hat{\mu}_{W}=0.70$). The distribution of the multiplicative errors was much more symmetric than that of the additive ones (Supplemental Figure S11). There were still some neighborhoods with much more extreme errors than others; interestingly, they differed from the neighborhoods with the worst additive ones. The most severe multiplicative error ($W=0.063$) was seen for Census Tract 315.04 within Randolph County, which was among the neighborhoods with additive errors under one mile in magnitude ($U=-0.92$). The standard deviation of the multiplicative errors $\hat{\sigma}_{W}=0.15$ was chosen as the moderate setting for simulations (Section 3.3).
@@ 4.5 Imputation models for straight-line proximity
For each health outcome, the imputation of map-based proximity to healthy foods $X$ began by fitting the linear regression imputation model described in Section 2.3. Specifically, $X$ was modeled conditionally on straight-line proximity $X^{*}$, metropolitan status $Z$, and the log-transformed outcome $\log(Y)$. An added interaction between $X^{*}$ and $Z$ allowed for different error mechanisms in metropolitan versus non-metropolitan neighborhoods, as is expected, for example, due to infrastructure differences. For diagnosed diabetes, $\hat{\mu}=-0.08+1.67X^{*}+0.88Z-0.42X^{*}\times Z-0.09\log(Y)\text{ and }\hat{\sigma}=0.96$. For obesity, $\hat{\mu}=1.11+1.67X^{*}+0.90Z-0.42X^{*}\times Z-0.24\log(Y)\text{ and }\hat{\sigma}=0.95$. For simplicity and given the high adjusted $R^{2}\approx 0.93$, other covariates (e.g., road density) were not included in the imputation models, but doing so would be straightforward.
Using these estimates, each neighborhood missing $X$ in the partially queried dataset was multiply imputed $B=20$ times for the imputation analyses. Imputing the missing $X$ values for the unqueried neighborhoods allowed them to contribute to the disease models in Section 4.6. With only minor modifications, missing $X$ values could also be imputed to supplement the subset of queried data and create a more accurate and complete description of the landscape of food access in the Piedmont Triad than the naive and complete case analyses, respectively. In Figure 1, the map of proximity to healthy foods based on a single deterministic imputation, not including $\log(Y)$ or resampling, is (i) far more comprehensive than the one based on the complete case (ignoring data for $339$ of $387$ neighborhoods) and (ii) more similar to the gold standard than the one based on the naive (using error-prone straight-line data).
@@ 4.6 Modeling the connection between proximity and disease
Finally, the connections between proximity to healthy foods and the prevalence of each health outcome across neighborhoods (census tracts) in the Piedmont Triad were modeled using Poisson regression. In contrast to the simulations, neighborhoods closer together in space could be more similar than neighborhoods farther away. To evaluate this potential spatial correlation, Moran’s I was calculated for both diseases of interest using the spdep package in R. [61] Specifically, Moran’s I was calculated for the residuals $e_{i}^{*}=Y_{i}-\widehat{Y}_{i}^{*}$ from the naive analysis, where $Y_{i}$ and $\widehat{Y}_{i}^{*}=\exp(\hat{\beta}_{0}^{*}+\hat{\beta}_{1}^{*}X_{i}+\hat{\beta}_{2}^{*}Z_{i}+\hat{\beta}_{3}^{*}X_{i}\times Z_{i})Pop_{i}$ denote the observed and predicted number of cases in neighborhood $i$, respectively. The parameters $(\hat{\beta}_{0}^{*},\hat{\beta}_{1}^{*},\hat{\beta}_{2}^{*},\hat{\beta}_{3}^{*})$ were taken from the naive models, which would have fully available data in practice and should be highly correlated with the gold standard ones. Both diagnosed diabetes ($I=0.06$, $p=0.02$) and obesity ($I=0.05$, $p=0.03$) had spatial autocorrelation that would be considered statistically significant at the $\alpha=0.05$ significance level.
@@
A conditional autoregressive (CAR) model was used to capture the associations between proximity to healthy foods and health outcomes (adjusting for metropolitan status) while allowing for spatial autocorrelation between neighborhoods that share a border. This model was defined as a mixed-effects Poisson regression with a random intercept for each neighborhood based on its bordering neighborhoods. Thus, instead of Equation (1),
$\log\{\textrm{E}_{{\boldsymbol{\beta}}}(Y|X,Z,r)\}=\beta_{0}+\beta_{1}X+\beta_{2}Z+\beta_{3}X\times Z+r+\log(Pop)$ (2)
@@
was of interest in the Piedmont Triad data, where $r$ denotes a neighborhood-specific random intercept. The vector of all neighborhoods’ random intercepts $\boldsymbol{r}=(r_{1},\dots,r_{397})^{\top}$ was assumed to follow a multivariate normal distribution with mean vector $\boldsymbol{0}$ and covariance matrix $\boldsymbol{\Sigma}$ defined to allow bordering neighborhoods to be correlated. See Waller and Gotway[62] for more technical details on the CAR model. While the random intercepts $\boldsymbol{r}$ could be correlated with each other, they were independent of the access variables $(X_{1},\dots,X_{397})^{\top}$ and metropolitan indicators $(Z_{1},\dots,Z_{397})^{\top}$. Using the fitme function from the spaMM package in R, [63] Equation (2) was estimated following the gold standard, naive, complete case, and imputation analysis approaches. For reference, the non-spatial models estimating Equation (1) were also fit to examine the impact of ignoring the spatial autocorrelation.
Additional simulations comparing the gold standard, naive, complete case, and imputation approaches for the mixed-effects model are included in Supplemental Section S.2. Under different amounts of variability in the spatial random effect, the imputation estimator exhibited low bias ($<2\%$), coverage probabilities close to the nominal $95\%$, and much higher relative efficiency to the gold standard ($0.64-0.67$) than the complete case analysis ($0.08-0.10$). More results from these simulations can be found in Supplemental Table S4.
@@ 4.7 Findings from the Piedmont Triad
Notice that metropolitan status $Z$ and its interaction with proximity $X$, were included in the analysis model in Equation (2). This inclusion modified the interpretation of our parameters and broadened our interest from just the coefficient on $X$ to include the coefficient on the interaction $X\times Z$. Now, $PR_{X}=\exp(\beta_{1})$ denotes the prevalence ratio between two non-metropolitan neighborhoods whose proximities to healthy foods differed by $1$ mile, and $PR_{XZ}=\exp(\beta_{3})$ denotes the ratio between the prevalence ratios for a 1-mile proximity difference in metropolitan versus non-metropolitan census neighborhoods. While $PR_{X}$ and $PR_{XZ}$ were of primary interest, the prevalence ratio $PR_{Z}=\exp(\beta_{2})$ between metropolitan and non-metropolitan tracts, adjusting for proximity, could also be helpful. Forest plots of the estimated adjusted prevalence ratios from each analysis method, along with their 95% confidence intervals (95% CI), are included in Figure 4.
@@
Figure 4: Estimated adjusted prevalence ratios (with 95% confidence intervals) for proximity to healthy foods, metropolitan status, and their interaction with the two health outcomes (diabetes and obesity) in the Piedmont Triad, North Carolina using four different analysis methods. Within each health outcome and method, estimates on the right with the dashed error bars came from the mixed-effects model allowing for spatial autocorrelation between bordering neighborhoods; estimates on the left with the solid error bars came from the non-spatial model assuming independence between neighborhoods.
4.7.1 Gold standard analysis
Based on the gold standard analysis, worse access to healthy foods (i.e., larger distances $X$ to the nearest healthy food store) was associated with a slightly lower prevalence of diagnosed diabetes in non-metropolitan neighborhoods ($\widehat{PR}_{X}=0.99$, 95% CI: $0.97,1.01$). That is, for every $1$ mile farther a neighborhood’s nearest healthy foods retailer was from its population center, its prevalence of diagnosed diabetes was expected to decrease by $1\%$. However, the confidence interval captured the null ($PR_{X}=1$). The same was found for metropolitan neighborhoods($\widehat{PR}_{XZ}=1.00$, 95% CI: $0.98,1.03$). Rather than seeing worse access (i.e., farther proximity) associated with a higher prevalence of diabetes, as was previously seen, [8, 10] our results were inconclusive about the direction of the effect. Importantly, our choice of proximity as an access measure, while those cited used density, captured a different facet of the food neighborhood environment. After adjusting for proximity, metropolitan neighborhoods were expected to have $18\%$ lower prevalence of diagnosed diabetes, on average ($\widehat{PR}_{Z}=0.82$, 95% CI: $0.74,0.91$).
For obesity, the estimated prevalence ratio for proximity to healthy foods among non-metropolitan census tracts from the gold standard analysis was at the null ($\widehat{PR}_{X}=1.00$, 95% CI: $0.99,1.01$). Estimates for metropolitan status ($\widehat{PR}_{Z}=0.99$, 95% CI: $0.94,1.04$) and its interaction with proximity ($\widehat{PR}_{XZ}=1.00$, 95% CI: $0.98,1.01$) were null, as well. The association between food access and obesity has seen conflict across the literature. Ahern et al. found that a higher density of some types of retailers (full-service restaurants and direct farm sales) led to lower rates, while other types (grocery stores) led to higher ones.[8] Meanwhile, others have found, like us, that there is not a relationship between them. [55]
4.7.2 Naive analysis
Despite the errors in straight-line proximity, the naive estimates of prevalence ratios for proximity to healthy foods on diagnosed diabetes ($\widehat{PR}_{X^{*}}=0.99$, 95% CI: $0.96,1.01$) and obesity ($\widehat{PR}_{X^{*}}=1.00$, 95% CI: $0.98,1.01$) in non-metropolitan neighborhoods closely resembled the gold standard and led to the same conclusions about an absence of associations. (We use the subscript $X^{*}$ here to denote the naive estimates based on $X^{*}$ rather than $X$.) Again, the diagnosed diabetes prevalence was expected to be lower in metropolitan neighborhoods ($\widehat{PR}_{Z}=0.82$, 95% CI: $0.74,0.91$), adjusting for proximity. In contrast, obesity prevalence was not expected to differ ($\widehat{PR}_{Z}=0.99$, 95% CI: $0.93,1.04$). Estimates on the interaction between metropolitan status and proximity to healthy foods were near the null for both outcomes (diagnosed diabetes [$\widehat{PR}_{X^{*}Z}=1.01$, 95% CI: $0.98,1.04$] and obesity [$\widehat{PR}_{X^{*}Z}=1.00$, 95% CI: $0.98,1.01$]).
4.7.3 Complete case analysis
@@
As expected, deleting the $339$ unqueried census tracts led to the complete case analysis having wider confidence intervals for all coefficients and outcome models. In addition, despite its use of map-based proximity measures, the point estimates differed from the gold standard analysis. For instance, the estimated prevalence ratios for proximity among non-metropolitan census tracts on diagnosed diabetes ($\widehat{PR}_{X}=0.98$, 95% CI: $0.96,1.01$) and obesity ($\widehat{PR}_{X}=0.99$, 95% CI: $0.98,1.00$) were both farther from the null. Notably, the complete case analysis should be asymptotically consistent since $X$ is missing at random (MAR) for the unqueried neighborhoods, [64] but there are no such guarantees in a small sample like this one ($n=48$). These differences could also be explained by the higher estimated baseline prevalence, i.e., the predicted disease prevalence in a non-metropolitan neighborhood where proximity to healthy foods was 00 miles (Supplemental Figure S12).
Again, the confidence intervals for the coefficient on the interaction captured the null for both outcomes (diagnosed diabetes [$\widehat{PR}_{XZ}=1.01$, 95% CI: $0.98,1.04$] and obesity [$\widehat{PR}_{XZ}=1.01$, 95% CI: $0.99,1.02$]), suggesting that the associations between proximity and disease prevalence were similar in metropolitan or non-metropolitan neighborhoods. Still, both estimates were farther from the null; for obesity, the estimate was even greater than one when the gold standard and naive estimates were less than or equal to it. Despite similar estimates on metropolitan status, the wider confidence interval led to a different conclusions about diagnosed diabetes ($\widehat{PR}_{Z}=0.87$, 95% CI: $0.73,1.03$) in the complete case than the gold standard and naive analyses, but the same conclusion about obesity ($\widehat{PR}_{Z}=0.99$, 95% CI: $0.90,1.09$).
4.7.4 Imputation analysis